# RaffeSDG: Random Frequency Filtering enabled Single-source Domain Generalization for Medical Image Segmentation

### v1

@@ Abstract
Deep learning models often encounter challenges in making accurate inferences when there are domain shifts between the source and target data. This issue is particularly pronounced in clinical settings due to the scarcity of annotated data resulting from the professional and private nature of medical data. Despite the existence of decent solutions, many of them are hindered in clinical settings due to limitations in data collection and computational complexity. To tackle domain shifts in data-scarce medical scenarios, we propose a Random frequency filtering enabled Single-source Domain Generalization algorithm (RaffeSDG), which promises robust out-of-domain inference with segmentation models trained on a single-source domain. A filter-based data augmentation strategy is first proposed to promote domain variability within a single-source domain by introducing variations in frequency space and blending homologous samples. Then Gaussian filter-based structural saliency is also leveraged to learn robust representations across augmented samples, further facilitating the training of generalizable segmentation models. To validate the effectiveness of RaffeSDG, we conducted extensive experiments involving out-of-domain inference on segmentation tasks for three human tissues imaged by four diverse modalities. Through thorough investigations and comparisons, compelling evidence was observed in these experiments, demonstrating the potential and generalizability of RaffeSDG. The code is available at https://github.com/liamheng/Non-IID_Medical_Image_Segmentation.
{IEEEkeywords}
@@ 1 Introduction
To facilitate the application of SDG, we introduce a Radom frequency filtering enabled SDG algorithm (RaffeSDG), which mitigates the computational and network training costs. RaffeSDG employs random frequency filtering and homologous sample blending to conveniently introduce randomization within the single-source domain through data augmentation. Furthermore, RaffeSDG also incorporates Gaussian filter-based structural saliency to construct a generalizable segmentor by learning robust representations in segmentation. Three human tissues imaged by four modalities are leveraged in the experiment to demonstrate the efficacy of RaffeSDG in out-of-domain inference. Our main contributions are summarised as follows:
@@
2 Related Work
2.1 Domain adaptation and domain generalization
@@
2.2 Single-source domain generalization
@@ 3.1 Definition and overview
3.2 Data augmentation for domain randomization
@@
3.2.1 Frequency filters based on Fourier transform
@@
The LFS variations can be introduced by adjusting the parameters of $D_{0}$ and $n$ (the parameter range is interpreted in Sec. 4.2). Notably, as depicted in Fig. 2 (a), distinct parameters are applied to the filter across different channels, promising the diversity of filtered samples.
@@
3.2.2 Homologous sample blending
@@
Consequently, the cooperation of random frequency filtering and homologous sample blending enables sub-image level data augmentation. According to Fig. 3 (c), this augmentation approach achieves remarkable in the single-source domain.
@@
3.3 Generalizable segmentor using structural saliency
@@ 4 Experiments
4.1 Experiment settings
Table 1: Experimental data summary
@@ 4.1 Experiment settings
4.1.1 Datasets
The experiments utilized four medical imaging modalities obtained from three different human tissues, as summarized in Table 1. The data for the vessel and retinal layer were publicly available, while the data for the joint cartilage were collected in collaboration with our partner, Southern University of Science and Technology Hospital, using discrepant ultrasound configurations of gain and grayscale. In order to generalize and assess segmentation models across these datasets, the ground-truth annotations were standardized. The datasets of DRIVE, HC-MS, and A were respectively employed as the single-source domain for each tissue, while the remaining datasets were treated as unseen target domains.
@@ 4.1.2 Implementation
4.2 Data augmentation configuration in RaffeSDG
@@
4.2.1 Filter types and blending strategies
To validate the configuration of the proposed data augmentation approach in RaffeSDG, we explored different filter types and blending strategies. The filtered and blended samples are illustrated in Fig. 4. Moreover, training data is augmented from DRIVE, and a vanilla U-Net is trained to quantitatively evaluate the performance of various configurations on the target domains, as summarized in Table 2.
@@ 4.2.1 Filter types and blending strategies
Filter type: To analyze the impact of filter types on data augmentation, we employ both the ideal filter and the Butterworth filter to introduce variations into the LFS of images. Fig.4 (b) presents the function diagram and the filtered result of the ideal filter. The sharp cutoff frequency of the ideal filter leads to the presence of ringing artifacts in the filtered result. On the other hand, the Butterworth filter offers a smoother transition between the passband and stopband, reducing the occurrence of ringing artifacts. Thus, the Butterworth filter is deemed more suitable for data augmentation, as illustrated in Fig.4 (c). Additionally, to enhance diversity in the augmented samples, we implement a channel-wise filter by applying Butterworth filters with different parameters to the RGB channels, as depicted in Fig. 4 (d). Consequently, superior performance is achieved by the channel-wise filter in Table 2.
Blending strategy: To interpret the effect of blending, we compare three strategies, including patch, grid, and continuous mask, as illustrated in Fig.4 (e). The images blended using patch and grid masks exhibit a noticeable gap, which deviates from real-world scenarios and hampers the overall context. In contrast, the continuous mask enables a seamless context in the augmented image, resulting in a more coherent representation that aligns well with real-world scenarios. The evaluation results provided in Table2 further validate our choice of filter and blending configurations. Accordingly, the channel-wise filter and the continuous mask are cooperated to implement RaffeSDG.
Table 2: Augmentation configuration comparison on training U-Net using only DRIVE for out-of-domain segmentation.
@@ 4.3 Augmentation comparison
The comparison includes both commonly used augmentation techniques and augmentation approaches designed in SDG algorithms. Specifically, the augmentation techniques include geometric transformation, CutOut [20], and CutMix [21]. And the SDG augmentation approaches are collected from ACVC [12], SLAug [13], GIN-IPA [14], and FreeSDG [15]. In the comparison, the dataset of DRIVE is utilized as the single-source domain, and a vanilla U-Net serves as the unified segmentor, which is trained with augmented data and conducts out-of-domain inference. The augmented samples are visualized in Fig. 7, while the segmentation performance is quantified in Table 3 by training a U-Net.
@@
As depicted in Fig. 7, the native augmentation strategy of geometric transformation, CutOut [20], and CutMix [21] are constrained to the single-source domain, leading to limited appearance variations in the augmented samples. In contrast, the augmentation approaches used in SDG tasks present remarkable performance in Table 3. For implementing SDG in classification, ACVC [12] introduces 19 visual corruptions and 3 Fourier-based corruptions to augment the single-source domain. For medical image segmentation, SLAug [13] introduces a saliency-balancing fusion strategy that combines global-scale and local location-scale aspects to augment the single-source domain. Moreover, GIN-IPA [14] proposes a causality-inspired data augmentation approach, which utilizes global intensity non-linear augmentation to diversify the appearances of medical images and addresses confounders through interventional pseudo-correlation augmentation.
Table 3: Augmentation technique comparison on training U-Net. The top three results are indicated in red, magenta, and blue.
@@
In our previous study, FreeSDG [15] utilized Gaussian filters to introduce LFS variations into samples, and combined with patch masks to merge the filtered results. However, although randomization was introduced through the augmentation of FreeSDG [15], the single-source domain, limited diversity was observed in Fig. 7 (f), resulting in suboptimal performance as indicated in Table 3. Fortunately, the augmentation approach proposed in RaffeSDG successfully introduces a wide range of visual variations in the single-source domain, as illustrated in Fig. 7 (h). This diversification strategy has proven to be effective, surpassing alternative methods in training the U-Net model and achieving superior performance, as demonstrated in Table 3.
Table 4: Comparison with SOTA segmentation algorithms. The top three results are indicated in red, magenta, and blue.
@@
4.4 Comparison with segmentation algorithms
Comparisons with SOTA segmentation and SDG algorithms are conducted to validate the advantages of the proposed RaffeSDG. The visualized results are presented in Fig. 6, while quantitative evaluations are summarized in Table 4 and 5.
4.4.1 Comparison with segmentation algorithms
@@
The comparative algorithms are trained based on their data dependency, with the DRIVE dataset serving as the default training data. The segmentation results are summarized in Table 4. Due to the lack of specific design for handling domain shifts, CS-Net [23] and UTNet [24] trained on the DRIVE dataset experience a notable decline in performance when applied to other datasets. By incorporating unlabeled test data during model training, DAMAN [6] and CSCADA [7] bridge domain shifts, resulting in improved performance in the target domains. Furthermore, FACT [9] and FedDG [10] utilize multiple source domains to learn generalizable segmentation models, demonstrating decent performance on unseen target domains.
@@
Table 5: Comparison with SDG algorithms.
@@
4.4.2 Comparison with SDG algorithms
Four SDG algorithms, including SAN-SAW [16], SLAug [13], GIN-IPA [14], and FreeSDG [15] are also compared and summarized in Table 5.
@@ 4.4.2 Comparison with SDG algorithms
4.5 Ablation studies
4.5.1 Data volume impact
@@
4.5.2 Ablation against segmentor modules
@@
Table 6: Comparisons of OCT and ultrasound segmentation. The top three results are indicated in red, magenta, and blue.
@@
4.6 Versatility for various tissues and imaging modalities
Comparisons of out-of-domain inference for OCT and ultrasound image segmentation were also performed to validate the effectiveness of RaffeSDG. Visualized and quantitative summaries of these comparisons are respectively presented in Fig. 9 and Table 6.
Considering that SAN-SAW [16] and GIN-IPA [14] employ specific designs for multi-category labels, it is understandable that they achieve remarkable performance in the retina layer segmentation of OCT, as expressed in Table 6. However, the nature of OCT and ultrasound data poses limitations for RaffeSDG. Unlike colorful fundus data obtained through visible light, OCT and ultrasound utilize wave propagation characteristics to capture images of internal structures, resulting in an inherently grayscale appearance. Consequently, OCT and ultrasound do not lend themselves well to the advantages of RaffeSDG, as illustrated in Fig. 9. Fortunately, despite these constraints, RaffeSDG still manages to achieve decent results, highlighting its versatility in medical image segmentation.
5 Discussions

### v2

@@ Abstract
Deep learning models often encounter challenges in making accurate inferences when there are domain shifts between the source and target data. This issue is particularly pronounced in clinical settings due to the scarcity of annotated data resulting from the professional and private nature of medical data. Despite the existence of decent solutions, many of them are hindered in clinical settings due to limitations in data collection and computational complexity. To tackle domain shifts in data-scarce medical scenarios, we propose a Random frequency filtering enabled Single-source Domain Generalization algorithm (RaffeSDG), which promises robust out-of-domain inference with segmentation models trained on a single-source domain. A frequency filter-based data augmentation strategy is first proposed to promote domain variability within a single-source domain by introducing variations in frequency space and blending homologous samples. Then Gaussian filter-based structural saliency is also leveraged to learn robust representations across augmented samples, further facilitating the training of generalizable segmentation models. To validate the effectiveness of RaffeSDG, we conducted extensive experiments involving out-of-domain inference on segmentation tasks for three human tissues imaged by four diverse modalities. Through thorough investigations and comparisons, compelling evidence was observed in these experiments, demonstrating the potential and generalizability of RaffeSDG. The code is available at https://github.com/liamheng/Non-IID_Medical_Image_Segmentation.
Index Terms:
@@ I Introduction
To facilitate the application of SDG, we introduce a Random frequency filtering enabled SDG algorithm (RaffeSDG), which mitigates the computational and network training costs. RaffeSDG employs random frequency filtering and homologous sample blending to conveniently introduce randomization within the single-source domain through data augmentation. Furthermore, RaffeSDG also incorporates Gaussian filter-based structural saliency to construct a generalizable segmentor by learning robust representations in segmentation. Three human tissues imaged by four modalities are leveraged in the experiment to demonstrate the efficacy of RaffeSDG in out-of-domain inference. Our main contributions are summarised as follows:
@@
II Related Work
II-A Domain adaptation and domain generalization
@@
II-B Single-source domain generalization
@@ III-A Definition and overview
III-B Data augmentation for domain randomization
@@
III-B1 Frequency filters based on Fourier transform
@@
The LFS variations can be introduced by adjusting the parameters of $D_{0}$ and $n$ (the parameter range is interpreted in Sec. IV-B). Notably, as depicted in Fig. 2 (a), distinct parameters are applied to the filter across different channels, promising the diversity of filtered samples.
@@
III-B2 Homologous sample blending
@@
Consequently, the cooperation of random frequency filtering and homologous sample blending enables sub-image level data augmentation. According to Fig. 3 (c), this augmentation approach achieves remarkable variability in the single-source domain.
@@
III-C Generalizable segmentor using structural saliency
@@ IV Experiments
IV-A Experiment settings
TABLE I: Experimental data summary
@@ IV-A Experiment settings
IV-A1 Datasets
The experiments utilized four medical imaging modalities obtained from three different human tissues, as summarized in Table I. The data for the vessel and retinal layer were publicly available, while the data for the joint cartilage were collected in collaboration with our partner, Southern University of Science and Technology Hospital, using discrepant ultrasound configurations of gain and grayscale. In order to generalize and assess segmentation models across these datasets, the ground-truth annotations were standardized. The datasets of DRIVE, HC-MS, and A were respectively employed as the single-source domain for each tissue, while the remaining datasets were treated as unseen target domains.
@@ IV-A2 Implementation
IV-B Data augmentation configuration in RaffeSDG
@@
IV-B1 Filter types and blending strategies
To validate the configuration of the proposed data augmentation approach in RaffeSDG, we explored different filter types and blending strategies. The filtered and blended samples are illustrated in Fig. 4. Moreover, training data is augmented from DRIVE, and a vanilla U-Net is trained to quantitatively evaluate the performance of various configurations on the target domains, as summarized in Table II.
@@ IV-B1 Filter types and blending strategies
Filter type: To analyze the impact of filter types on data augmentation, we employ both the ideal filter and the Butterworth filter to introduce variations into the LFS of images. Fig.4 (b) presents the function diagram and the filtered result of the ideal filter. The sharp cutoff frequency of the ideal filter leads to the presence of ringing artifacts in the filtered result. On the other hand, the Butterworth filter offers a smoother transition between the passband and stopband, reducing the occurrence of ringing artifacts. Thus, the Butterworth filter is deemed more suitable for data augmentation, as illustrated in Fig.4 (c). Additionally, to enhance diversity in the augmented samples, we implement a channel-wise filter by applying Butterworth filters with different parameters to the RGB channels, as depicted in Fig. 4 (d). Consequently, superior performance is achieved by the channel-wise filter in Table II.
Blending strategy: To interpret the effect of blending, we compare three strategies, including patch, grid, and continuous mask, as illustrated in Fig.4 (e). The images blended using patch and grid masks exhibit a noticeable gap, which deviates from real-world scenarios and hampers the overall context. In contrast, the continuous mask enables a seamless context in the augmented image, resulting in a more coherent representation that aligns well with real-world scenarios. The evaluation results provided in TableII further validate our choice of filter and blending configurations. Accordingly, the channel-wise filter and the continuous mask are cooperated to implement RaffeSDG.
TABLE II: Augmentation configuration comparison on training U-Net using only DRIVE for out-of-domain segmentation.
@@ IV-C Augmentation comparison
The comparison includes both commonly used augmentation techniques and augmentation approaches designed in SDG algorithms. Specifically, the augmentation techniques include geometric transformation, CutOut [20], and CutMix [21]. And the SDG augmentation approaches are collected from ACVC [12], SLAug [13], GIN-IPA [14], and FreeSDG [15]. In the comparison, the dataset of DRIVE is utilized as the single-source domain, and a vanilla U-Net serves as the unified segmentor, which is trained with augmented data and conducts out-of-domain inference. The augmented samples are visualized in Fig. 7, while the segmentation performance is quantified in Table III by training a U-Net.
@@
As depicted in Fig. 7, the native augmentation strategy of geometric transformation, CutOut [20], and CutMix [21] are constrained to the single-source domain, leading to limited appearance variations in the augmented samples. In contrast, the augmentation approaches used in SDG tasks present remarkable performance in Table III. For implementing SDG in classification, ACVC [12] introduces 19 visual corruptions and 3 Fourier-based corruptions to augment the single-source domain. For medical image segmentation, SLAug [13] introduces a saliency-balancing fusion strategy that combines global-scale and local location-scale aspects to augment the single-source domain. Moreover, GIN-IPA [14] proposes a causality-inspired data augmentation approach, which utilizes global intensity non-linear augmentation to diversify the appearances of medical images and addresses confounders through interventional pseudo-correlation augmentation.
TABLE III: Augmentation technique comparison on training U-Net. The top three results are indicated in red, magenta, and blue.
@@
In our previous study, FreeSDG [15] utilized Gaussian filters to introduce LFS variations into samples, and combined with patch masks to merge the filtered results. However, although randomization was introduced through the augmentation of FreeSDG [15], the single-source domain, limited diversity was observed in Fig. 7 (f), resulting in suboptimal performance as indicated in Table III. Fortunately, the augmentation approach proposed in RaffeSDG successfully introduces a wide range of visual variations in the single-source domain, as illustrated in Fig. 7 (h). This diversification strategy has proven to be effective, surpassing alternative methods in training the U-Net model and achieving superior performance, as demonstrated in Table III.
TABLE IV: Comparison with SOTA segmentation algorithms. The top three results are indicated in red, magenta, and blue.
@@
IV-D Comparison with segmentation algorithms
Comparisons with SOTA segmentation and SDG algorithms are conducted to validate the advantages of the proposed RaffeSDG. The visualized results are presented in Fig. 6, while quantitative evaluations are summarized in Table IV and V.
IV-D1 Comparison with segmentation algorithms
@@
The comparative algorithms are trained based on their data dependency, with the DRIVE dataset serving as the default training data. The segmentation results are summarized in Table IV. Due to the lack of specific design for handling domain shifts, CS-Net [23] and UTNet [24] trained on the DRIVE dataset experience a notable decline in performance when applied to other datasets. By incorporating unlabeled test data during model training, DAMAN [6] and CSCADA [7] bridge domain shifts, resulting in improved performance in the target domains. Furthermore, FACT [9] and FedDG [10] utilize multiple source domains to learn generalizable segmentation models, demonstrating decent performance on unseen target domains.
@@
TABLE V: Comparison with SDG algorithms.
@@
IV-D2 Comparison with SDG algorithms
Four SDG algorithms, including SAN-SAW [16], SLAug [13], GIN-IPA [14], and FreeSDG [15] are also compared and summarized in Table V.
@@ IV-D2 Comparison with SDG algorithms
IV-E Ablation studies
IV-E1 Data volume impact
@@
IV-E2 Ablation against segmentor modules
@@
TABLE VI: Comparisons of OCT and ultrasound segmentation. The top three results are indicated in red, magenta, and blue.
@@
IV-F Versatility for various tissues and imaging modalities
Comparisons of out-of-domain inference for OCT and ultrasound image segmentation were also performed to validate the effectiveness of RaffeSDG. Visualized and quantitative summaries of these comparisons are respectively presented in Fig. 9 and Table VI.
Considering that SAN-SAW [16] and GIN-IPA [14] employ specific designs for multi-category labels, it is understandable that they achieve remarkable performance in the retina layer segmentation of OCT, as expressed in Table VI. However, the nature of OCT and ultrasound data poses limitations for RaffeSDG. Unlike colorful fundus data obtained through visible light, OCT and ultrasound utilize wave propagation characteristics to capture images of internal structures, resulting in an inherently grayscale appearance. Consequently, OCT and ultrasound do not lend themselves well to the advantages of RaffeSDG, as illustrated in Fig. 9. Fortunately, despite these constraints, RaffeSDG still manages to achieve decent results, highlighting its versatility in medical image segmentation.
V Discussions